OpenStudy (anonymous):

The Ratio of roses to snapdragons was 4 to 5. If there were 26,000 roses on the float, how many snapdragons were there? PLEASE HELP ME :C AND PLEASE SHOW YOUR WORK !

OpenStudy (anonymous):

okay,do 26,000/4

OpenStudy (anonymous):

did you get 6,500

OpenStudy (anonymous):

now multiply 6,500 by 5

OpenStudy (anonymous):

to get 32,500

OpenStudy (anonymous):

then u add them ?

OpenStudy (anonymous):

no, the ratio is 26,000: 32,500

OpenStudy (anonymous):

so 32,500 snapdragons

OpenStudy (anonymous):

oh , i get it thanks , so much your so helpful c:

OpenStudy (anonymous):

wait , one second

OpenStudy (anonymous):

i need to check if I did it right

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

I am not entirely sure

OpenStudy (anonymous):

I can tag some other people just in case

OpenStudy (anonymous):

@Skielerlucas04

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

let's see.. we can re-write the ratios as fractions so given roses to snapdragons was 4 to 5. 4 roses: 5 snapdragons now we are given 26,000 roses and ? snapdragons so that's the unknown we need to solve for \[\large \frac{4}{5}=\frac{26,000}{x}\] then we can use cross multiplication \[\large 4(x) = 26,000(5)\] \[\large 4x = 130,000\] dividing 4 on both sides gives me \[\large x=32,500\] snapdragons so our ratio is 26,000 roses to 32,500 snapdragons or 26,000:32,500 yes @heretohelpalways you've done it right.

OpenStudy (anonymous):

@UsukiDoll Thank you :)